Circular No. 6577

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

GRB 970228 D. Palmer, T. Cline, N. Gehrels, P. Kurczynski, R. Ramaty, H. Seifert,

and B. Teegarden, Goddard Space Flight Center, NASA; and K. Hurley, University of California at Berkeley, report: “TGRS data have been analyzed for the most intense 3.6 s of this γ-ray burst, corresponding to the initial peak mentioned on IAUC 6572. The spectrum is consistent with that of a classical burst. The peak countrate in the band 40–200 keV is 3.3 counts s^{−1} cm^{−2} (1.2-s average). This low countrate precludes tight constraints on spectral form and parameters, although a lack of detectable flux above 300 keV indicates a high-energy cutoff steeper than 10^{−3}. An optically-thin thermal bremsstrahlung spectral form with temperature 100 keV < T < 150 keV fits the data adequately over the interval 35–1000 keV. This spectral fit gives a fluence of 2 × 10^{−6} erg/cm^{2} over this time interval (although a comparable fluence may exist in the remainder of the burst). Assuming that the x-ray emission seen by BeppoSAX (IAUC 6576) has a 1-day exponential decay time, the persistent 0.5–10-keV emission has a total energy fluence that is ∼ 10–20 percent of the prompt γ-ray fluence.”

I. A. Smith and E. P. Liang, Rice University; and R. A. Gruendl and K. Y. Lo, University of Illinois at Urbana/Champaign, report: “Observations at 86.4 GHz (3.5 mm) using the Berkeley-Illinois-Maryland Association array were made on Mar. 7 UT, centered on the VLA source reported by Frail et al. (IAUC 6576). No significant source was detected there or anywhere else inside the x-ray error box of SAX J0501.7+1146 (Costa et al., IAUC 6576) with a 1σ rms of 1.2 mJy.”

B. Margon and E. W. Deutsch, University of Washington; and J. Secker, Washington State University, write: “Observations with the ARC 3.5-m telescope on Mar. 7 and 8 UT reveal a faint (R = 21) image coincident (± 1′′) with the VLA position of the radio source (Frail et al., IAUC 6576) that lies in the BeppoSAX Narrow Field position (Costa et al., IAUC 6576) for the x-ray source possibly associated with GRB 970228 (IAUC 6572). If the x-ray and optical sources are identical, which of course remains to be proven, then the x-ray flux observed on Mar. 4 (IAUC 6576) implies an xray-to-optical flux ratio consistent with that of an active galactic nucleus; however, that ratio in the brighter x-ray state of Feb. 28, closer in time to the γ-ray burst, would be extraordinary.”

1997 March 8  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green